OpenStudy (anonymous):

Which is the set of roots for the equation? x2 + 3x - 8 = 0 a: {18,-1} b: {-6,3} c: {-2,9} d: {9}

OpenStudy (campbell_st):

I think the equation should be \[x^2 + 3x - 18 = 0\] is that correct..?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

yeah it should

OpenStudy (anonymous):

Which is the set of roots for the equation? x2 + 3x - 18 = 0 a: {18,-1} b: {-6,3} c: {-2,9} d: {9}

OpenStudy (anonymous):

what do you guys think

OpenStudy (jdoe0001):

hint: \(\large \begin{array}{cccllll} x^2&+3x&-18\\ &\uparrow &\uparrow \\ &6-3&6\cdot -3 \end{array}\)

OpenStudy (anonymous):

so b

OpenStudy (jdoe0001):

well.. try those 2 roots in "b", see if they work

OpenStudy (jdoe0001):

have you covered factoring of quadratics yet?